Perform the indicated operation and simplify your answer.
(5x - 3)(4x - 7)

Answer by bcampbell23669(14)

20x^2-47x+21
The foil method is used to solve this
Firsts
Outters
Inners
Lasts
5x(4x) = 20x^2
5x(-7) = -35x
-3(4x) = -12x
-3(-7) = 21
-35x (+) -12x = -47x
Answer: 20x^2-47x+21
Hope this helps. Good luck :-)